OpenStudy (anonymous):

The lateral area of a cone is 500mc012-1.jpg cm2. The radius is 50 cm. Find the slant height to the nearest tenth. 10 cm 9.1 cm 14.2 cm 11.2 cm

OpenStudy (anonymous):

\[500pi cm ^{2}\]

OpenStudy (anonymous):

\[A_{l} = \Pi \times R \times S\] where: Al - lateral area R - radius S - slant height and \[S = \frac{ A_{l} }{ \Pi \times R } = \frac{ 500\Pi }{ \Pi \times 50 } = 10\]